OpenStudy (anonymous):

Write the standard form of the equation of the line passing through the point (1,3) and perpendicular to the line x-2y=-8.

OpenStudy (ybarrap):

You need an equation of the form $$ y=mx+b $$ The line perpendicular to x-2y=-8 will have a slope that is the negative inverse of this line: $$ x-2y=-8\\ -2y=-8-x\\ y=(1/2)x+4 $$ Since the slope of this equation is (1/2), the slope of the line perpendicular to it will have slope -2: $$ y=-2x+b $$ Since the lines must intersect at point (1,3), set x=1 and y=3 to find b for your perpendicular equation: $$ 3=-2(1)+b\\ b=3+2=5 $$ so the line perpendicular to x-2y=-8 and going through point (1,3) is $$ y=-2x+5 $$ Make sense?